Find an equation in standard form of the line passing through the point (2,5) with slope m=5
Answer by ankor@dixie-net.com(22740)

Find an equation in standard form of the line passing through the point (2,5) with slope m = 5.
:
Use the point/slope formula: y - y1 = m(x - x1)
:
given: m = 5; x1 = 2; y1 = 5, substitute:
:
y - 5 = 5(x - 2)
:
y - 5 = 5x - 10
:
y = 5x - 10 + 5
:
y = 5x - 5
:
The standard form: ax + by = c
:
-5x + y = -5; subtracted 5x from both sides
or
+5x - y = +5; multiplied eq by -1
: